OpenStudy (staceyg):

Write y=(- 3/4)x+3 in standard form using integers.

OpenStudy (anonymous):

multiply by 4 both sides and arrnge the terms such that right hand side has only 0

OpenStudy (staceyg):

3x + 4y=12 3x + 4y= -12 3x - 4y = 12 -3x + 4y = 12

OpenStudy (staceyg):

Do you know the formula?

OpenStudy (anonymous):

see 4*y =4*((- 3/4)x+3 ) now simplify

OpenStudy (anonymous):

4y = -3x +12 now add 3x to both sides we have 3x +4y =-3x+3x +12 3x+4y =12

OpenStudy (staceyg):

Thanks :) So for y= -0.5x + 0.8. How would you do that?

OpenStudy (anonymous):

y= -(1/10) x + 8/10 so 10y = -x +8 or x+10y = 8

OpenStudy (staceyg):

not -x+10y=8?

OpenStudy (staceyg):

@matricked

OpenStudy (anonymous):

NO

OpenStudy (staceyg):

lol alright thanks :)